# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

### v1

@@ Abstract
Reward baseline is important for Reinforcement Learning (RL) algorithms to reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) is proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to inaccurate advantage estimates in environments with highly noisy rewards, potentially introducing bias. In this work, we propose a model, called Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), by using lightweight Kalman filtering to dynamically estimate the latent reward mean and variance. This filtering technique replaces the naive batch mean baseline, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through experiments and analyses, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 1 Introduction
In reinforcement learning algorithms that directly optimize the policy via gradient estimates, including Policy Gradient (PG) (sutton1999policy, ), Actor-Critic (AC) (mnih2016asynchronous, ) and Proximal Policy Optimization (PPO) (schulman2017proximal, ), variance reduction of policy gradient estimate is critical for stable and sample-efficient training. A common approach to estimate the advantage of actions is to subtract a baseline, which can be interpreted as the expectation of returns by taking all actions for the current state, from the observed returns. By doing so, the model ideally should approximate the expected value of the return under the current policy. In many popular algorithms, such as AC and PPO, this baseline is often computed via an additionally maintained value model. But it will introduce further computational cost, especially for large model training.
To overcome this issue, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) is proposed to simplify the prediction of reward baselines by taking the mean reward over a batch or group of trajectories. While this approach can introduce bias if inaccurate action advantage was obtained when the reward distribution is highly non-stationary and sparse. This bias can lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we take one step ahead by treating the observed rewards as noisy measurements of a latent reward signal and use the Kalman filter to model the underlying rewards and noises. The resulting filtered baseline adapts over time, smoothing out noise while tracking underlying values and uncertainties in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with existing GRPO algorithm and requires no modification to the policy networks. It provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach, and show it improves training stability.
In summary, this paper makes the following contributions:
@@
  • We empirically demonstrate the proposed KRPO algorithm can receive higher rewards than its the standard group-mean counterparts across different datasets.
@@ 2.1 Deep Reinforcement Learning
Among model-free approaches, value-based methods learn to approximate the expected return of actions under a given policy. The Deep Q-Network (DQN) (mnih2013playing, ) is a foundational algorithm that learns a Q-function using temporal difference learning following Bellman Equation. Several extensions improve its stability and performance, including Double DQN (van2016deep, ) to address Q-value overestimation; Dueling DQN (wang2016dueling, ), which decomposes Q-values into state-value and advantage terms.
@@ 2.2 Reinforcement Learning to Improve Reasoning for LLMs
Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks underlying uncertainties. This gets better advantage estimation with minimal computational cost.
@@ 3.1 Problem Setting
We consider reinforcement learning where an agent interacts with an environment modeled as a Markov Decision Process (MDP). At a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantage function to reduce the variance of gradient updates.
In Group Relative Policy Optimization (GRPO), this framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting this mean. While this reduces variance, the group mean may be an inaccurate estimate of the expected reward for individual outputs, particularly when the reward distribution is noisy. This can lead to unstable or suboptimal updates during training.
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose to replace the naive group mean baseline with a Kalman filter that maintains an online estimate of the reward distribution’s mean and variance. Each observed return $r_{i}$ is treated as a noisy measurement of the latent reward signal $x_{i}$.
The Kalman filter maintains a posterior estimate of the reward mean $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of $x_{i}$ given observations up to step $i-1$, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $Q$ is the process noise variance and $\sigma^{2}$ is the measurement noise variance. These are hyperparameters that control the smoothness and adaptability of the estimated baseline.
@@ 3.3 KRPO Algorithm
By take the posterior variance estimation $P_{i|i}$ into account, the advantage can be more effective to learn. That is to say, the “good” actions with large rewards can be better enhanced and the “poor” actions with small rewards can be further punished.
@@
Hyperparameters: KL weight $\beta$, group size $n$, batch size $B$, learning rate $\eta$, clipping $\epsilon$, process noise $Q$, measurement noise $\sigma^{2}$ Initialize replay buffer $\mathcal{B}$, optimizer $\text{Adam}(\theta,\eta)$ ;
@@ 4 Connections of Reward Mean and Kalman Filtering
This empirical mean can be interpreted as a special case of a Kalman filter with the following assumptions: (1) The process noise variance $Q\to 0$, indicating that the latent state (true reward mean) is assumed to be constant; (2) The measurement noise variance $\sigma^{2}$ is a constant; (3) no prior information for the initial state.
Specifically, consider the Kalman filter update equations as aforementioned Eq. 1 to Eq. 5. When $\sigma^{2}$ is constant and the system state does not change over time (i.e., the state transition matrix is the identity and the process noise is zero), the Kalman filter update reduces to a weighted average of the observations, where the weights are determined by the prior variance and the observation noise variance. In practice, however, when the full batch of $\{r_{i}\}$ is available, computing the sample mean corresponds to performing a batch update using all data equally. This is equivalent to a Bayesian update with all measurements weighted equally, which can be seen as a limiting form of the Kalman filter without temporal dynamics.
This observation motivates the use of a Kalman filter, which allows $Q>0$, and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive and sequential-aware estimate of the reward baseline.
@@ 5.1 Data
We use two datasets in our experiments, i.e. Arithmetic Dataset and OpenMathInstruct Dataset, each of which is divided into three sub-datasets: easy, normal, and hard.
@@ 5.2 Implementation Details
We adopt group size of 12 for both GRPO and our proposed KRPO models. Training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation managable, we adopt Llama-3.2-1B-Instruct as our base model, to individually trained with GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and variance of the filter measurement noise $\sigma^{2}$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, the training is conducted for 5,500 steps. For the reward design, following DeepSeek-R1-Zero model (guo2025deepseek, ) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48 Gigabyte of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.
@@ 5.3.1 Overall Performance of KRPO
Table 1: Model accuracy comparison between the KRPO with other models on Arithmetic and OpenMath-Instruct datasets. The best performance for each block with the same setting are bolded.
@@
We compare the accuracy performance of our proposed Kalman-filter-based Relative Policy Optimization (KRPO) against several baselines, including the original Llama3.2 model, Proximal Policy Optimization (PPO), and Group Relative Policy Optimization (GRPO). We calculated the accuracy by also assigning half correctness if the ground-truth answer is contained in the returned strings, e.g. ground-truth is “59”, but “The answer is 59” is considered half correct. Table 1 shows the results on two mathematical reasoning benchmarks: Arithmetic and OpenMath-Instruct. The results are reported across three difficulty levels: Easy, Normal, and Hard.
@@
Beyond the overall performance improvements, we observe interesting trends across the two datasets. We find that the raw pretrained Llama3.2 model achieves higher accuracy on Arithmetic-Easy (22.516%) compared to OpenMath-Easy (13.056%), likely due to the short and highly patterned nature of Arithmetic questions, which allows the language model to leverage surface-level cues. However, after applying reinforcement learning-based reasoning training, KRPO model trained on OpenMath outperform the accuracy on Arithmetic. This highlights that Arithmetic problems, although syntactically simpler, actually require strong numerical reasoning ability, which is the weak point of Llama3.2 models. In contrast, OpenMath problems are longer and more linguistically complex, but once the reasoning structure is learned, the model can generalize more effectively across examples.
@@ 5.3.2 Training Reward Curve Comparison between KRPO and GRPO
Figure 1: The curves of returns/rewards within a batch for different difficulty levels of questions within Arithmetic Dataset.
@@
We also show the training return curves comparison between KRPO and GRPO on Arithmetic Dataset as Fig. 1 and on OpenMath-Instruct Dataset as Fig. 2. We use running average for reward smoothing to have a better visualization quality. We set the group size to 12 and rollout batchsize to 32. In this case, the maximal reward can be obtained at each point is 384.
On both datasets and across all difficulties, the proposed KRPO algorithm can consistently converge faster and eventually reach much higher rewards. Interestingly, on the OpenMath-Instruct dataset, the performance gap between KRPO and GRPO increases with problem difficulty. This trend suggests that KRPO provides more stable and accurate advantage estimates in harder reasoning scenarios, where variance in reward signals is likely higher and a naive baseline such as mean reward becomes less effective. This trend is also applied for the Arithmetic easy and normal subset, but not observed on the Arithmetic hard dataset. It is possibly caused by the Arithmetic hard data distribution difference with other difficulty sets.
@@ 5.4.1 KL Loss Weights $\alpha$
We visualize the training curves with different KL weights in the loss $\alpha$ of KRPO by setting it to {0, 0.001, 0.01, 0.05} as Fig. 3(a). We observe that with the 0.01 $\alpha$ reaches the fastest convergence, as well as the highest returns. However, with 0.001 settings, the rewards are converged slower but eventually reaches the same point; with 0.05 and 0 settings the reward are much worse than 0.01. So $\alpha=0.01$ is set as the default setting for KRPO.
@@ 5.4.2 Model Convergence against Process Noise and Measurement Noise
We investigate the model convergence against process noise $Q$ and measurement noise $\sigma^{2}$ as Fig. 4(b). We noticed that by changing the process noise $Q$ from 1e-5 to 1e-3 and changing the measurement noise $\sigma^{2}$ from 1e-2 to 1e-1 does not change the final convergence points, just making the convergence slower. However, if make the measurement noise $\sigma^{2}$ smaller to 1e-3, the model performance will be drastically worse.

### v2

@@ Abstract
Reward baseline is important for Reinforcement Learning (RL) algorithms to reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) is proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to inaccurate advantage estimates in environments with highly noisy rewards, potentially introducing bias. In this work, we propose a model, called Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), by using lightweight Kalman filtering to dynamically estimate the latent reward mean and variance. This filtering technique replaces the naive batch mean baseline, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through accuracy and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 1 Introduction
In reinforcement learning algorithms that directly optimize the policy via gradient estimates, including Policy Gradient (PG) (sutton1999policy, ), Actor-Critic (AC) (mnih2016asynchronous, ) and Proximal Policy Optimization (PPO) (schulman2017proximal, ), variance reduction of policy gradient estimate is critical for stable and sample-efficient training. A common approach to estimate the advantage of actions is to subtract a baseline, which can be interpreted as the expectation of returns by taking all actions for the current state, from the observed returns. By doing so, the model is ideally expected to approximate the expected value of the return under the current policy. In many popular algorithms, such as AC and PPO, this baseline is often computed via an additionally maintained value model. However, the value model introduces further computational costs, especially for large model training.
To overcome this issue, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) is proposed to simplify the prediction of reward baselines by taking the mean reward over a batch or group of trajectories. But this approach can introduce bias if inaccurate action advantage was estimated when the reward is highly non-stationary and sparse. This bias may lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we take one step ahead by treating the observed rewards as noisy measurements of a latent reward signal and use the Kalman filter to model the underlying reward and noise trajectories. The resulting filtered baseline adapts over time, smoothing out noise while tracking underlying values and uncertainties in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with the existing GRPO algorithm and requires no modification to the policy networks. It provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach and show how it improves training stability. In summary, this paper makes the following contributions:
@@
  • We empirically demonstrate that the proposed KRPO algorithm can receive higher rewards than its standard group-mean counterparts across different datasets.
@@ 2.1 Deep Reinforcement Learning
Among model-free approaches, value-based methods learn to approximate the expected return of actions under a given policy. The Deep Q-Network (DQN) (mnih2013playing, ) is a foundational algorithm that learns a Q-function using temporal difference learning following the Bellman Equation. Several extensions improve its stability and performance, including Double DQN (van2016deep, ) to address Q-value overestimation; Dueling DQN (wang2016dueling, ), which decomposes Q-values into state-value and advantage terms.
@@ 2.2 Reinforcement Learning to Improve Reasoning for LLMs
Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks underlying uncertainties. With non-parametric one-dimensional Kalman filter, we can have more accurate advantage estimation by filtering out the underlying reward baseline and uncertainty with minimal computational cost.
@@ 3.1 Problem Setting
We consider reinforcement learning where an agent interacts with an environment modeled as a Markov Decision Process (MDP). At a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, and receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantage function to reduce the variance of gradient updates.
Group Relative Policy Optimization (GRPO) framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting the group sample mean. While this reduces variance, the group mean may be an inaccurate estimate of the expected reward for individual outputs, particularly when the reward distribution is noisy. This can lead to unstable or suboptimal updates during training.
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose to replace the naive group mean baseline with a Kalman filter that maintains an online estimate of the reward distribution’s mean and variance. Each observed return $r_{i}$ is treated as a noisy measurement of the latent reward signal $x_{i}$ which we cannot obtain directly.
The Kalman filter maintains a posterior estimate of the reward baseline $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of $x_{i}$ given observations up to step $i-1$, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $Q$ is the process noise variance and $R$ is the measurement noise variance. These are hyperparameters that control the smoothness and adaptability of the estimated baseline.
@@ 3.3 KRPO Algorithm
By taking the posterior variance estimation $P_{i|i}$ into account, the advantage can be more effective to learn. That is to say, the “good” actions with large rewards can be better enhanced and the “poor” actions with small rewards can be further punished.
@@
Hyperparameters: KL weight $\beta$, group size $n$, batch size $B$, learning rate $\eta$, clipping $\epsilon$, process noise $Q$, measurement noise $R$ Initialize replay buffer $\mathcal{B}$, optimizer $\text{Adam}(\theta,\eta)$ ;
@@ 4 Connections of Reward Mean and Kalman Filtering
This empirical mean can be interpreted as a special case of a Kalman filter with the following assumptions: (1) The process noise variance $Q\to 0$, indicating that the latent state (expected reward) is assumed to be constant; (2) The measurement noise variance $R$ is a constant; (3) no prior information for the initial state.
Specifically, consider the Kalman filter update equations as aforementioned Eq. 1 to Eq. 5. When $Q$ is constant and the system state does not change over time (i.e., the state transition matrix is an identity matrix and the process noise is zero), the Kalman filter update reduces to a weighted average of the observations, where the weights are determined by the prior variance and the observation noise variance. In practice, however, when the full batch of $\{r_{i}\}$ is available, computing the sample mean corresponds to performing a batch update using all data equally. This is equivalent to a Bayesian update with all measurements weighted equally, which can be seen as a limiting form of the Kalman filter without temporal dynamics.
This observation motivates the use of a Kalman filter, which allows slight noises when new rewards observed ($Q>0$), and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive and sequential-aware estimate of the reward baseline.
@@ 5.1 Data
We use two datasets in our experiments, i.e., the Arithmetic Dataset and OpenMathInstruct Dataset, each of which is divided into three sub-datasets: easy, normal, and hard.
@@ 5.2 Implementation Details
We adopt a group size of 12 for both GRPO and our proposed KRPO models. The training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation managable, we adopt Llama-3.2-1B-Instruct as our base model, to individually train with GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and variance of the filter measurement noise $R$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, the training is conducted for 5,500 steps. For the reward design, following DeepSeek-R1-Zero model (guo2025deepseek, ) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48 Gigabyte of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.
@@ 5.3.1 Overall Performance of KRPO
Table 1: Model accuracy comparison between the KRPO and other models on the Arithmetic and OpenMath-Instruct datasets. The best performance for each block with the same setting are bolded.
@@
We compare the accuracy performance of our proposed Kalman-filter-based Relative Policy Optimization (KRPO) against several baselines, including the original Llama3.2 model, Proximal Policy Optimization (PPO), and Group Relative Policy Optimization (GRPO). We calculated the accuracy by also assigning half correctness if the ground-truth answer is contained in the returned strings, e.g., ground-truth is “59”, but “The answer is 59” is considered half correct. Table 1 shows the results on two mathematical reasoning benchmarks: Arithmetic and OpenMath-Instruct. The results are reported across three difficulty levels: Easy, Normal, and Hard.
@@
Beyond the overall performance improvements, we observe interesting trends across the two datasets. We find that the raw pretrained Llama3.2 model achieves higher accuracy on Arithmetic-Easy (22.516%) compared to OpenMath-Easy (13.056%), likely due to the short and highly patterned nature of Arithmetic questions, which allows the language model to leverage surface-level cues. However, after applying reinforcement learning-based reasoning training, the KRPO model trained on OpenMath outperforms the accuracy on Arithmetic. This highlights that Arithmetic problems, although syntactically simpler, actually require strong numerical reasoning ability, which is the weak point of Llama3.2 models. In contrast, OpenMath problems are longer and more linguistically complex, but once the reasoning structure is learned, the model can generalize more effectively across examples.
@@ 5.3.2 Training Reward Curve Comparison between KRPO and GRPO
Figure 1: The curves of returns/rewards within a batch for different difficulty levels of questions within the Arithmetic Dataset.
@@
We also show the training return curves comparison between KRPO and GRPO on the Arithmetic Dataset as Fig. 1 and on the OpenMath-Instruct Dataset as Fig. 2. We use a running average for reward smoothing to improve visualization quality. We set the group size to 12 and the rollout batch size to 32. In this case, the maximal reward that can be obtained at each point is 384.
On both datasets and across all difficulties, the proposed KRPO algorithm can consistently converge faster and eventually reach much higher rewards. Interestingly, on the OpenMath-Instruct dataset, the performance gap between KRPO and GRPO increases with problem difficulty. This trend suggests that KRPO provides more stable and accurate advantage estimates in harder reasoning scenarios, where variance in reward signals is likely higher and a naive baseline such as mean reward becomes less effective. This trend is also applied to the Arithmetic easy and normal subset, but not observed on the Arithmetic hard dataset. It is possibly caused by the Arithmetic hard data distribution difference with other difficulty sets.
@@ 5.4.1 KL Loss Weights $\alpha$
We visualize the training curves with different KL weights in the loss $\alpha$ of KRPO by setting it to {0, 0.001, 0.01, 0.05} as Fig. 3(a). We observe that with the 0.01 $\alpha$, it reaches the fastest convergence, as well as the highest returns. However, with 0.001 settings, the rewards are converged more slowly but eventually reach the same point; with 0.05 and 0 settings, the rewards are much worse than 0.01. So $\alpha=0.01$ is set as the default setting for KRPO.
@@ 5.4.2 Model Convergence against Process Noise and Measurement Noise
We investigate the model convergence against process noise $Q$ and measurement noise $R$ as Fig. 4(b). We noticed that by changing the process noise $Q$ from 1e-5 to 1e-3 and changing the measurement noise $R$ from 1e-2 to 1e-1 does not change the final convergence points, just making the convergence slower. However, if we make the measurement noise $R$ smaller to 1e-3, the model performance will be drastically worse.